# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v2

@@ 6.1 Measure-over-Measure Comparison
We validate SURF against ICE-Eval and C-SHAP-Eval. As there is no “groundtruth” of faithfulness to compare evaluation results to, we test the frameworks in three manufactured settings, where we know the relative faithfulness of the explanations a priori. Specifically, we look at faithfulness results for a perfect explanation and two randomized explanations. To generate the Perfect explanation, observe that the most faithful explanation of the model is the model itself. Thus, we set $\{A_{i},V_{i}\}_{i=1}^{C}$ according to the weights of the linear classification layer. In our second setting (termed Rand Imp), we keep the perfect CAVs fixed and randomly sample the concept importances. Our final setting (termed Full Rand) additionally randomly samples the CAVs. In the Perfect setting, we expect to obtain no faithfulness error, regardless of the evaluation measure. Then, we expect to obtain increasingly less faithful scores as we progressively increase randomness (from Rand Imp to Full Rand).
@@
We evaluate the faithfulness in these three settings on an ImageNet-pretrained ResNet-50 [16] finetuned on the Caltech-101 dataset [21]. Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq. 6 and ICE’s normalized L1 logit error (Norm L1) defined in Eq. 5. As a complementary measure to EMD, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs. This measure ensures that fine-grained changes that result in a different prediction ordering are penalized but ignores differences in magnitudes. Finally, we measure surrogate complexity with the number of learnable parameters the evaluation measure introduces. Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics. More details in Appendix Sec. A.
@@
Results are reported in Tab. 3, averaged over 10 seeds. In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected. However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations. In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because they do not use concept importances, so even random concept importances have the same result as the Perfect setting. Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics. Finally, in the Full Rand setting, C-SHAP-Eval unintuitively reports an improvement in faithfulness compared to the Perfect setting. ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected. Of the tested faithfulness measures, SURF is the only one that behaves as expected, as validated by all metrics.
@@
Next, we turn our attention to the metrics. We highlight a failure case of Top-1 by looking at C-SHAP-Eval (CEL)’s results in the Full Rand setting. The Top-1 score is strong; however, the other metrics (especially EMD) reveal that the reconstructed output has severe errors on the remainder of the output distribution. Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1), despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits. We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across other classes (as measured by MAE) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
As shown in the three settings presented, SURF is able to appropriately measure the faithfulness of three toy U-CBEMs. This contrasts with other approaches and metrics, which fail to yield satisfactory measurements in the same three cases. The remainder of the paper uses SURF to evaluate prior U-CBEMs.
@@ 6.2 Benchmarking U-CBEMs
We evaluate five prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, and MCD). We benchmark U-CBEMs on a varied range of tasks with a diverse range of models to demonstrate the applicability of SURF. Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression. In all tasks, each U-CBEM discovers 5 CAVs (or subspaces for MCD) per output. CAVs and importances are found on the training set, and the resulting explanations are evaluated on the test set. Dataset, finetuning, and U-CBEM implementation details are in Sec. B. Good-faith efforts were made to adapt prior U-CBEMs (that were developed for object classification) to new tasks.
@@
We evaluate on three tasks. (1) Object Classification: We finetune a ResNet-50 on Caltech-101 and report MAE, EMD, Top-1, and Rank Corr. (2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [31] to predict the presence/absence of attributes in the CelebA dataset [22]. We report MAE and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1. (3) Age Estimation: Given a human face, we finetune a ViT [6] to predict their age (in years) on the UTK-Face dataset [45]. Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used. C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs. For this task, we report MAE.
@@
Benchmark results are reported in Tab. 4. Aside from MCD, we observe that no prior U-CBEM is faithful to the original model, on any task. In the Object Classification task, the most faithful U-CBEM (after MCD) achieves the lowest EMD of 0.438, denoting large errors in the probability-space which stem from errors in the logit-space. Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task. CDISCO exhibits an average error of 32.6 years in the Age Estimation task.
@@
We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding. Crucially, the class-specific CAVs are found on images from only that class, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will result in large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors. This issue is exacerbated for CRAFT & ICE, which non-linearly project to the concept space. (2) Except for ICE and MCD, the concept importance scores do not satisfy a logit-based completeness criterion (note, C-SHAP satisfies an accuracy-based completeness criterion). Thus, the concept importances do not faithfully reflect the impact of the concept representation on the model’s outputs. Though MCD also finds class-specific concepts, MCD achieves superior faithfulness by allowing a concept to be the span of many CAVs (a linear subspace); thus, it is able to faithfully explain a larger part of the model’s reasoning with the same number of concepts.
@@
Figure 2: We fit U-CBEMs on Sec. 6.2’s Object Classification task with an increasing number of concepts and evaluate them with SURF. Some U-CBEMs do not improve as the number of concepts increase. U-CBEMs that improve quickly plateau, indicating an “elbow” phenomenon.
@@
The most important hyperparameter for any U-CBEM is the number of concepts to be discovered. Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation. However, presenting users with explanations containing too many concepts may be uninterpretable. The cognitive psychology literature has shown that humans can hold a limited number of items in their working memory at a time [5, 25]. Thus, simple, or parsimonious, explanations are preferred, which can communicate the bulk of the model’s reasoning in a few concepts. Instead of setting this hyperparameter arbitrarily, something commonly done in prior U-CBEMs (10 in [44], 25 in [9, 13]), we use SURF to intelligently analyze this tradeoff.
Using the Object Classification task from Sec. 6.2, we fit U-CBEMs with an increasing number of concepts. Each U-CBEM is evaluated using the SURF measure, visualized in Fig. 2. We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts. Only ICE and MCD uniformly improve as the number of concepts increases, and they exhibit an “elbow” phenomenon, marking a natural choice for the number of concepts.
Table 5: Novel setting: faithfulness on adversarial images. U-CBEMs are fixed from Sec. 6.2’s Object Classification task. The original model is adversarially attacked, and we report the mean paired difference in MAE and EMD (adversary - original) on 100 randomly sampled images. Any statistically insignificant difference is omitted (with “-”). U-CBEMs are generally not robust to adversarial attack; all of them exhibit an increase in MAE.
U-CBEM MAE ($\downarrow$) EMD ($\downarrow$)
CDISCO ++0.634 ++0.039
ICE ++0.730 -
CRAFT ++0.411 ++0.073
C-SHAP ++0.631 -0.195
MCD ++0.339 -0.038
6.4 Faithfulness in Adversarial Settings
Prior work has restricted their faithfulness evaluations to explanations on images from the same dataset the U-CBEM was trained on, as we did in Sec. 6.2. Such evaluations do not test U-CBEMs in challenging situations; yet, we believe these situations are the ones where having faithful explanations are the most critical. One such scenario is adversarial attack, which we test here. We take on the role of an adversary who has white-box access to the original model, but is unaware of the post-hoc explanation. If the U-CBEM is robust to adversarial attack, then we expect the U-CBEM to faithfully explain the model when it is under adversarial attack, allowing stakeholders to be aware of the adversarial attack. We design a Projected Gradient Descent (PGD) attack [24] to maximize the probability distribution difference of the original model on the input image and the adversary (details in Appendix Sec. C). The PGD attack is repeated on a 100 randomly sampled images from the training set for U-CBEMs fit in the Object Classification task from Sec. 6.2.
Tab. 5 reports the mean paired difference in MAE and EMD (adversarial - original). We use a one-sided, paired t-test on per-image differences to determine statistically significant increases in error (if $p<0.01$). We find that the MAE error increase is significant for all U-CBEMs, while EMD only significantly increases for CDISCO and CRAFT. We conclude that all U-CBEMs are less robust in the logit space, with ICE being the most vulnerable. The logit-space errors only propagate to the probability space for certain U-CBEMs (CDISCO & CRAFT). In certain cases (e.g., C-SHAP), the probability-space error may actually decrease on adversarial examples. We anticipate that U-CBEMs may be more vulnerable to adversarial attacks specifically designed to fool the explainer (instead of the model); as demonstrated under a simple PGD attack, SURF makes it possible to study such scenarios.
@@ 7 Conclusion
This paper organizes prior faithfulness measures, discusses their limitations, and proposes SURF to accurately measure faithfulness. We validate SURF compared to other measures in a measure-over-measure study, and we use SURF to find that several existing U-CBEMs produce unfaithful explanations. We apply SURF to downstream settings (number of concepts selection and adversarial robustness) to demonstrate its utility. We urge future works to adopt SURF to comprehensively evaluate faithfulness and compare to prior works.
@@
    Sec. A: Implementation details for the surrogate justification experiment (Sec. 6.1).
@@
  3. 3.
    Sec. C: Experimental details for the adversarial attack experiment done in Sec. 6.4.
@@ A.2 Surrogate Implementation
Recall that U-CBEMs define their own mechanism $\mathcal{P}:\mathcal{H}\to\mathbb{R}^{K}$ to project embeddings to the concept basis. For these experiments, since we are not using any U-CBEM and instead directly setting the CAVs and importances, we define $\mathcal{P}(\mathbf{h}_{j},V_{i})_{k}=\mathbf{h}_{j}^{T}\mathbf{v}_{i,k}$, where $i$ denotes the class, $k$ denotes the CAV, and $j$ selects the spatial embedding. Recall in this case that $K=1$ and there is only one embedding (obtained from global pooling).
A.2.1 SF Surrogate (Ours)
Following Sec. 5.1 and Eq. 10, we set the parameters of the surrogate with the concepts and importances obtained from the setting. This surrogate has no learnable parameters and is exactly equal to the original linear. Note that the embeddings are linearly projected onto the CAVs to find the concept representation.
@@ A.2.3 C-SHAP Surrogate
Since C-SHAP’s surrogate contains learnable parameters. Following [39], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities. We train the surrogate using the Adam optimizer [19] and a starting learning rate of $0.1$. The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus. We stop training when the learning rate is reduced below $10^{-7}$.
@@ B.1 Model Details
Object Classification: Our evaluation uses a ResNet-50. We use the ImageNet pre-trained weights and initialize the final classification layer with random weights. The model is finetuned (conv4 and all subsequent layers) on the Caltech-101 training set, achieving a test accuracy of approximately 92.75%. The model is trained without data augmentation and using cross entropy loss.

### v3

@@ 6.1 Measure-over-Measure Comparison
The fundamental challenge in evaluating faithfulness measures is that there is no “groundtruth” of faithfulness to compare evaluation results to. To address this, we propose a simple sanity check, which introduces three manufactured settings where the relative faithfulness of explanations is known a priori. Then, we check if the faithfulness measure preserves the relative ordering of scores across settings.
@@
Concretely, we look at faithfulness results for a perfect explanation and two randomized explanations. To generate the Perfect explanation, observe that the most faithful explanation of the model is the model itself. Thus, we set $\{A_{i},V_{i}\}_{i=1}^{C}$ according to the weights of the linear classification layer. In our second setting (termed Rand Imp), we keep the perfect CAVs fixed and randomly sample the concept importances. Our final setting (termed Full Rand) additionally randomly samples the CAVs. In the Perfect setting, we expect to obtain no faithfulness error, regardless of the evaluation measure. Then, we expect to obtain increasingly less faithful scores as we progressively increase randomness (from Rand Imp to Full Rand).
@@
We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [resnet] finetuned on the Caltech-101 dataset [caltech-101]. Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq. 6 and ICE’s normalized L1 logit error (Norm L1) defined in Eq. 5. As a complementary measure to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs. Rank Corr penalizes fine-grained changes that result in a different prediction ordering but ignores differences in magnitudes. Finally, we measure surrogate complexity by its learnable parameters and FLOPs needed to obtain the output from the concept representation. Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics. More details are in Appendix Sec. A.
@@
Results are reported in Tab. 3, averaged over 10 seeds. In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected. However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations. In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because their surrogates do not use concept importances, so using random concept importances has the same evaluation result as the Perfect setting. Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics. Finally, in the Full Rand setting, C-SHAP-Eval unintuitively reports an improvement in faithfulness compared to the Perfect setting. ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected.
@@
Next, we turn our attention to the metrics. We highlight a failure case of Top-1 by looking at SURF’s results in the Rand Imp setting. The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the reconstructed output has severe errors on the remainder of the output distribution. Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits. We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across all classes (as measured by $\text{SURF}_{\text{MAE}}$) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
Of the faithfulness measures, only SURF passes the sanity check while having the lowest surrogate complexity (fewest FLOPs and no learnable parameters). The remainder of the paper uses SURF to study prior U-CBEMs.
@@ 6.2 Benchmarking U-CBEMs
We evaluate six prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD). We benchmark U-CBEMs on a varied range of tasks and models to demonstrate the applicability of SURF. Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression. In all tasks, U-CBEMs discover 5 CAVs (or subspaces for MCD and HU-MCD) per output. CAVs and importances are found on the training set; the resulting explanations are evaluated on the test set. Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Sec. B.
@@
We evaluate on three tasks. (1) Object Classification: We use the Caltech-101-finetuned ResNet-50 and report $\text{SURF}_{\text{MAE}}$, $\text{SURF}_{\text{EMD}}$, Top-1, and Rank Corr. HU-MCD requires architectural modifications, so we can only evaluate it on this task, which uses a compatible ResNet-50. (2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [Sandler2018MobileNetV2IR] to predict the presence/absence of attributes on the CelebA dataset [celeba]. We report $\text{SURF}_{\text{MAE}}$ and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1. (3) Age Estimation: Given a human face, we finetune a ViT [VIT] to predict their age (in years) on the UTK-Face dataset [utkface]. Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used. C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs.
@@
Benchmark results are reported in Tab. 4. In the evaluated settings, we observe that no prior U-CBEM is faithful to the original model. In the Object Classification task, the most faithful U-CBEM is HU-MCD with an $\text{SURF}_{\text{EMD}}$ of 0.384, denoting large errors in the probability-space. Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task. CDISCO exhibits an average error of 32.6 years in the Age Estimation task.
@@
We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding. Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors. This issue is exacerbated for CRAFT and ICE, which non-linearly project to the concept space. Though MCD and HU-MCD also find class-specific concepts, they achieve superior faithfulness by using concept subspaces; thus they faithfully explain more of the model with the same number of concepts. (2) Most methods do not satisfy a logit-based completeness criterion, preventing a simple reproduction of model outputs. Note that C-SHAP satisfies an accuracy-based completeness criterion and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be presented to the user in the explanation.
@@
Figure 2: We fit U-CBEMs on Sec. 6.2’s Object Classification task with an increasing number of concepts and evaluate their faithfulness with SURF. Some U-CBEMs do not improve as the number of concepts increase. U-CBEMs that improve quickly plateau.
@@
The most important hyperparameter for any U-CBEM is the number of concepts to be discovered. Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation. However, explanations containing too many concepts may be difficult to interpret. Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [2010TheMM, Miller1956TheMN]. Thus, simple, or parsimonious, explanations are preferred, which communicate the bulk of the model’s computation with a few concepts. However, inadequate faithfulness measures make the faithfulness-parsimony tradeoff hard to analyze, so prior U-CBEMs commonly set this hyperparameter arbitrarily (10 in ICE, 25 in CRAFT & ACE). SURF allows us to intelligently analyze this tradeoff.
Using the Object Classification task from Sec. 6.2, we fit U-CBEMs with an increasing number of concepts. Each U-CBEM is evaluated using the SURF measure, visualized in Fig. 2. We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts. ICE marginally improves as we add more concepts. Only MCD and HU-MCD uniformly improve as the number of concepts increases, and they exhibit a plateauing effect, marking a natural choice for the number of concepts. Interestingly, HU-MCD is more faithful than MCD only when both discover relatively few concepts; as the number of concepts grows, MCD becomes more faithful. This offers a more nuanced perspective on the findings from the HU-MCD paper, where they claimed superior faithfulness over MCD based on results from their deletion-based proxy.
@@ 7 Conclusion
This paper argues that we lack a clear view on how faithful current U-CBEMs really are, largely because of measures that deviate from the formal definition of faithfulness. In accordance with this definition, we develop desiderata for a faithfulness measure and propose SURF as a simple, principled measure for faithfulness. SURF passes a basic sanity check, reveals that state-of-the-art U-CBEMs are substantially unfaithful, and is a practical tool for studying the tradeoff between parsimony and faithfulness when choosing the number of concepts. We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores alongside their interpretability claims.
[…]
@@
    Sec. A: Implementation details for the measure-over-measure study (Sec. 6.1).
@@ A.2 Surrogate Implementation
Recall that U-CBEMs define their own mechanism $\mathcal{P}:\mathcal{H}\to\mathbb{R}^{K}$ to project embeddings to the concept basis. For these experiments, since we are not using any U-CBEM and instead directly setting the CAVs and importances, we define $\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}=\mathbf{h}_{j}^{T}\mathbf{v}_{i,k}$, where $i$ denotes the class, $k$ denotes the CAV, and $j$ selects the spatial embedding. Recall in this case that $K=1$ and there is only one embedding (obtained from global pooling).
A.2.1 SURF Surrogate (Ours)
Following Sec. 5.1 and Eq. 10, we set the parameters of the surrogate with the obtained concepts and importances. This surrogate has no learnable parameters and is exactly equal to the original linear layer (up to the bias term). Note that the embeddings are linearly projected onto the CAVs to find the concept representation. After the surrogate reconstructs the output, we add the constant bias term before evaluation.
@@ A.2.3 C-SHAP Surrogate
Since C-SHAP’s surrogate contains learnable parameters. Following [ConceptSHAP], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities. We train the surrogate using the Adam optimizer [kingma2014adam] and a starting learning rate of $0.1$. The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus. We stop training when the learning rate is reduced below $10^{-7}$.
A.3 FLOPs Computation
We treat multiply and add as 2 FLOPs. Let $K$ be the number of concepts, $C$ be the number of output classes, and $D$ by the dimensionality of the embedding space.
SURF Surrogate: Given that the surrogate is a linear layer from concept representations to the output space (though it uses different concept representations for each class), the number of FLOPs required is $2KC$.
ICE Surrogate: This surrogate reconstructs the original embedding (by a linear projection) from the concept representation, and then uses the original model (which is also a linear projection) to map to the output space. Thus, the number of FLOPs required is $CD(2K+1)$.
C-SHAP Surrogate: This surrogate reconstructs the original embedding (with a two-layer MLP) from the concept representation, and then uses the original model (which is a linear projection) to map to the output space. Thus, the number of FLOPs required is $2C(H_{1}K+H_{1}D+D)$, where $H_{1}$ is the dimensionality of the MLP’s hidden layer.
@@ B.1 Model Details
Object Classification: Our evaluation uses a ResNet-50, available from the timm library. We use the ImageNet pre-trained weights and initialize the final classification layer with random weights. The model is finetuned (conv4 and all subsequent layers) on the Caltech-101 training set, achieving a test accuracy of approximately 95.61%. The model is trained without data augmentation and using cross entropy loss with label smoothing.